OpenStudy (anonymous):

solve: (3xe^x)+(x^2e^x)

OpenStudy (anonymous):

its not an equation

OpenStudy (anonymous):

\[3xe^x+x^2e^x=0\]

OpenStudy (anonymous):

The solutions are 0 and -3.

OpenStudy (anonymous):

First multiply by e^(-x) You may do that because e^x is never 0.

OpenStudy (anonymous):

Then you get 3x+x^2=x(3+x)